# SAC: Neural Speech Codec with Semantic-Acoustic Dual-Stream Quantization

### v1

@@ 4.1 Training Details
To ensure diversity in the training data, we randomly sampled approximately 20,000 hours of bilingual (Chinese and English) speech data from various sources. These include Emilia (He et al., 2024), WenetSpeech4TTS (Ma et al., 2024), LibriSpeech (Panayotov et al., 2015), Libriheavy (Kang et al., 2024), MLS (Pratap et al., 2020), and in-house data. All audio is resampled to 16 kHz. Further details of the training data are provided in Appendix A.
@@
Both the semantic and acoustic codebooks in SAC contain 16,384 entries. To provide configurations for different bitrates, we set the acoustic token frame rate to 25 Hz or 50 Hz, corresponding to overall token rates of 37.5 Hz and 62.5 Hz, respectively. Models are trained for 850k steps on 8 NVIDIA H20 GPUs with a batch size of 24. During training, each audio sample is randomly cropped into 2.4-second segments. Optimization is performed using the AdamW optimizer (Loshchilov and Hutter, 2017) with $\beta_{1}=0.8$ and $\beta_{2}=0.9$. Both the generator and discriminator learning rates are initialized at $1\times 10^{-4}$ and decayed exponentially throughout training. Additional training details are provided in Appendix B.
@@ 4.2 Evaluation Details
We evaluate speech reconstruction performance on the LibriSpeech test-clean set, which contains 2,620 utterances sampled at 16 kHz. To facilitate comparison across models, we report key parameters, including codebook size, the number of quantizers (Nq), token rate, and bandwidth (BPS). Speech intelligibility is assessed using Short-Time Objective Intelligibility (STOI) and Word Error Rate (WER), where transcriptions are obtained with the HuBERT-based (Hsu et al., 2021) model [^2] https://huggingface.co/facebook/hubert-large-ls960-ft. Acoustic quality is measured by the Perceptual Evaluation of Speech Quality (PESQ) and UTMOS (Saeki et al., 2022), while speaker similarity (SIM) is computed via a WavLM-based (Chen et al., 2022b) speaker verification model  [^3] https://github.com/microsoft/UniSpeech/tree/main/downstreams/speaker_verification. For comparison, we evaluate SAC at two token rates against a range of state-of-the-art codecs with similar bitrates, including DAC (Kumar et al., 2023), Encodec (Défossez et al., 2022), Mimi (Défossez et al., 2024), SpeechTokenizer (Zhang et al., 2023), SemantiCodec (Liu et al., 2024), BigCodec (Xin et al., 2024), the X-Codec series (Ye et al., 2025a, b), XY-Tokenizer (Gong et al., 2025), WavTokenizer (Ji et al., 2024b), MagiCodec (Song et al., 2025), and TS3-Codec (Wu et al., 2024). All baselines are reproduced on our test set using their official checkpoints. For ablations, we consistently adopt the lower-bitrate configuration of SAC for comparison.
@@
We evaluate the semantic richness of codec representations using the speech domain of the ARCH benchmark, following Ji et al. (2024b) and Jiang et al. (2025). ARCH includes RAVDESS (Livingstone and Russo, 2018) and EMOVO (Costantini et al., 2014) for emotion recognition, SLURP (Bastianelli et al., 2020) for intent classification, and AudioMNIST (Becker et al., 2024) for digit recognition. For each codec model, the quantized representations extracted from the codec quantizer are average-pooled over time and passed through a linear classifier, following the standard ARCH protocol. For semantic–acoustic decoupled codecs such as SemantiCodec and SAC, features from the two streams are concatenated before linear probing. To further contextualize the performance gap between discrete codec representations and continuous SSL features, we additionally report results for wav2vec 2.0 (Baevski et al., 2020), data2vec (Baevski et al., 2022), HuBERT (Hsu et al., 2021), and WavLM (Chen et al., 2022a) in ARCH as SSL-based references.
@@ 5.1 Speech Reconstruction Results
To further examine robustness, we additionally evaluate SAC’s reconstruction performance under noisy conditions, as presented in Appendix C.
@@ 5.2 Representations Evaluation Results
5.3 Effect of Auxiliary Feature Supervision
@@ 5.4 Speech Decoupling Analysis
As illustrated in Fig. 3, we performed a visual analysis of SAC’s reconstruction capabilities across different paradigms. In the full reconstruction setting, which follows SAC’s basic reconstruction method using both semantic and acoustic tokens, the spectrogram preserves more high-frequency textures compared to the original, indicating reduced distortion in harmonic structures and formants. This observation also explains SAC’s strong performance on UTMOS, as the decoder effectively acts as a generator to enrich fine-grained acoustic detail.
In contrast, the semantic-only reconstruction lacks the speaker-specific fundamental frequencies and formants, yet retains clear semantic content. This demonstrates SAC’s ability to effectively separate and preserve the semantic features, while discarding speaker-dependent acoustic information. On the other hand, the acoustic-only reconstruction retains well-defined fundamental frequencies and formants, but completely loses semantic content (The output is perceptually similar to noise, devoid of any meaningful information). This highlights SAC’s remarkable capacity to disentangle and independently process semantic and acoustic features, providing direct evidence of its robust disentanglement ability. Additional mel-spectrogram reconstruction analyses of other codec models are provided in Appendix D.
@@ 6 Conclusion
In this paper, we presented SAC, a neural speech codec with semantic–acoustic dual-stream quantization. To exploit the strengths of semantic tokenizers in capturing linguistic content and codecs in modeling acoustic detail, SAC introduces independent semantic and acoustic streams that extract semantic and acoustic tokens separately. To further enhance timbre modeling, we incorporated speaker feature supervision into codec training. Comprehensive evaluations demonstrate that SAC achieves SOTA performance in both speech reconstruction and semantic representation across different bitrates, while ablation studies confirm the effectiveness of auxiliary feature supervision. Furthermore, we observed a perfect disentanglement effect in SAC’s semantic and acoustic tokens during reconstruction. Specifically, the reconstruction based on semantic tokens completely lacks speaker-related information, while the reconstruction based on acoustic tokens retains no semantic content. To the best of our knowledge, this is the first instance of such a clean disentanglement between semantics and speaker identity in terms of reconstruction. This finding offers new insights for subsequent tasks that rely on disentanglement, such as voice conversion or the joint control of style and timbre in TTS.
@@
Although SAC demonstrates superior reconstruction quality and semantic representation within the speech domain, its generalizability to other types of audio signals, such as music and sounds, remains to be explored. A key challenge lies in the semantic tokenizer used in SAC, which is trained under ASR supervision on speech data and therefore primarily aligned with textual objectives. In contrast, the semantics in music and sound extend beyond linguistic alignment. To develop a more general-purpose audio codec based on the SAC dual-stream architecture, future work should focus on designing a universal audio semantic tokenizer.
@@ Appendix A Details of Training Data
As summarized in Table 6, SAC is trained on approximately 20,000 hours of bilingual (Chinese and English) speech, comprising around 8 million utterances. The training corpus includes roughly 10,000 hours in each language.
For Chinese, the data sources include the Chinese subset of Emilia (He et al., 2024), WenetSpeech4TTS (Ma et al., 2024), and a small amount of in-house para-linguistic data. For English, we use LibriSpeech (Panayotov et al., 2015), the small and medium subsets of Libriheavy (Kang et al., 2024), the English subset of Emilia, MLS (Pratap et al., 2020), and a small amount of in-house data.
To promote diversity, we follow the annotation scheme of VoxBox (Wang et al., 2025b), which labels speech samples along three dimensions: age, gender, and emotion. Specifically, we partition Emilia, WenetSpeech4TTS, and MLS into categories defined by the unique combination of these three attributes, and then sample data to achieve balanced coverage across categories. This strategy ensures broad diversity within the training set.
@@ Appendix B Codec Training Details
To accelerate SAC training, we pre-extracted all semantic representations $\mathbf{S}_{c}$ and corresponding semantic tokens from the training corpus using a pretrained semantic tokenizer (Zeng et al., 2024b). During training, only the codebook from the tokenizer is required to obtain the quantized semantic embeddings $\mathbf{S}_{q}$, while $\mathbf{S}_{c}$ serves as the ground-truth target for auxiliary semantic feature supervision. The generator contains approximately 277M parameters, with around 249M being trainable. The codebook of the semantic tokenizer and the speaker encoder are frozen to ensure that well-pretrained semantic and speaker features are properly utilized.
For generator training, the loss coefficients are set as $\lambda_{\text{recon}}=15$, $\lambda_{\text{vq}}=1$, $\lambda_{\text{adv}}=1$, $\lambda_{\text{feat}}=2$, $\lambda_{\text{sem}}=1000$, and $\lambda_{\text{spk}}=10$. Within the VQ loss, the commitment and codebook loss weights are set to 0.25 and 4, respectively. To improve training stability, the generator is pretrained for 1,500 steps before introducing the discriminator for adversarial training. An exponential moving average (EMA) is applied to maintain a smoothed version of the model parameters (Chen et al., 2025), which are used during inference and observed to enhance model stability.
@@ Appendix C Speech Reconstruction under Noisy Conditions
Notably, SAC maintains a clear advantage in UTMOS, SIM, and WER. For perceptual quality, SAC achieves UTMOS scores of 3.84 and 3.90 at high and low bitrates, respectively—both substantially higher than the ground truth value of 3.50. We attribute this to SAC’s acoustic stream effectively modeling fine-grained acoustic details, while the decoder functions as a generator that enhances perceptual realism. In terms of speaker similarity, SAC shows only a marginal drop of 0.01 in SIM compared to the clean condition, whereas other codecs experience significant declines. This demonstrates the robustness of both our model architecture and the diverse training data, as well as the effectiveness of the speaker feature supervision in preserving timbre characteristics. Furthermore, SAC maintains high speech intelligibility under noisy conditions, with remarkably low WER values, confirming that its dual-stream architecture enables reliable semantic reconstruction even in the presence of noise.
@@
Appendix D Mel-Spectrogram Reconstructions Comparison among Codecs

### v2

@@ 4.1 Training Details
To ensure diversity in the training data, we randomly sampled approximately 20,000 hours of bilingual (Chinese and English) speech data from various sources. These include Emilia (emilia), WenetSpeech4TTS (wenetspeech4tts), LibriSpeech (librispeech), Libriheavy (libriheavy), MLS (mls), and in-house data. Further details of the training data are provided in Appendix A.
@@
Both the semantic and acoustic codebooks in SAC contain 16,384 entries. To provide configurations for different bitrates, we set the acoustic token frame rate to 25 Hz or 50 Hz, corresponding to overall token rates of 37.5 Hz and 62.5 Hz, respectively. Models are trained for 850k steps on 8 NVIDIA H20 GPUs with a batch size of 24. During training, each audio sample is randomly cropped into 2.4-second segments. Optimization is performed using the AdamW optimizer (adamw) with $\beta_{1}=0.8$ and $\beta_{2}=0.9$. Both the generator and discriminator learning rates are initialized at $1\times 10^{-4}$ and decayed exponentially throughout training. Additional training details are provided in Appendix B.
@@ 4.2 Evaluation Details
We evaluate speech reconstruction performance on the LibriSpeech test-clean set, which contains 2,620 utterances at 16 kHz. To facilitate comparison across models, we report key parameters, including codebook size, the number of quantizers (Nq), token rate, and bandwidth (BPS). Speech intelligibility is assessed using Short-Time Objective Intelligibility (STOI) and Word Error Rate (WER), where transcriptions are obtained with the HuBERT-based (hubert) model. Acoustic quality is measured by the Perceptual Evaluation of Speech Quality (PESQ) and UTMOS (utmos), while speaker similarity (SIM) is computed via a WavLM-based (wavlm-sv) speaker verification model. For comparison, we evaluate SAC at two token rates against a range of state-of-the-art codecs with similar bitrates, including DAC (dac), Encodec (encodec), Mimi (moshi), SpeechTokenizer (speechtokenizer), SemantiCodec (semanticodec), BigCodec (bigcodec), the X-Codec series (xcodec; llasa-xcodec2), XY-Tokenizer (XY-Tokenizer), WavTokenizer (wavtokenizer), MagiCodec (magicodec), and TS3-Codec (Ts3-codec). All baselines are reproduced on our test set using their official checkpoints. For ablations, we consistently adopt the lower-bitrate configuration of SAC for comparison.
@@
We evaluate the semantic richness of codec representations using the speech domain of the ARCH benchmark, following wavtokenizer and unicodec. ARCH includes RAVDESS (RAVDESS) and EMOVO (EMOVO) for emotion recognition, SLURP (SLURP) for intent classification, and AudioMNIST (Audiomnist) for digit recognition. For each codec model, the quantized representations extracted from the codec quantizer are average-pooled over time and passed through a linear classifier, following the standard ARCH protocol. For semantic–acoustic decoupled codecs such as SemantiCodec and SAC, features from the two streams are concatenated before linear probing. To further contextualize the performance gap between discrete codec representations and continuous SSL features, we additionally report results for wav2vec 2.0 (wav2vec2), data2vec (data2vec), HuBERT (hubert), and WavLM (wavlm) in ARCH as SSL-based references.
@@ 5.1 Speech Reconstruction Results
We include a Mean Opinion Score (MOS) subjective evaluation to assess the perceptual quality of the reconstructed audio, with details in Appendix E. To further examine robustness, we also evaluate SAC’s reconstruction performance under noisy conditions, as presented in Appendix C.
@@ 5.2 Representations Evaluation Results
5.3 Effect of Auxiliary Feature Supervision
The impact of $\mathcal{L}_{\text{spk}}$ and $\mathcal{L}_{\text{sem}}$ on semantic representation is further analyzed in Appendix H.
@@ 5.3 Effect of Auxiliary Feature Supervision
5.4 Downstream Speech-LLM Performance
We conduct downstream LLM-based experiments to assess the transferability of SAC when integrated with speech language models. Given SAC’s strong semantic representations and its previously verified ASR performance (glm-4-voice), our focus here is on its effectiveness in generative tasks. We adopt a single-stage autoregressive (AR) TTS framework using the pre-trained Qwen3-0.6B (qwen3) as the backbone, and accommodate SAC’s dual-stream tokens through an interleaved flattening scheme based on their token-rate ratio. Using the 37.5 Hz SAC tokenizer, we train the TTS model on a 100k-hour bilingual (Chinese–English) corpus. As shown in Table 9, the resulting system substantially outperforms state-of-the-art pure AR models, including Spark-TTS (spark-tts) and Llasa (llasa-xcodec2), achieving significantly lower WER and higher UTMOS, which reflects superior semantic clarity and objective perceptual quality. Although speaker similarity exhibits a slight drop, we attribute this to the low-bitrate SAC variant used under current computational constraints. Detailed setup and analyses are provided in Appendix D.
@@ 5.5 Speech Decoupling Analysis
As illustrated in Fig. 3, we performed a visual analysis of SAC’s reconstruction capabilities across different paradigms. In the full reconstruction setting, which follows SAC’s basic reconstruction method using dual-stream tokens, the spectrogram preserves more high-frequency textures compared to the original, indicating reduced distortion in harmonic structures and formants. This observation also explains SAC’s strong performance on UTMOS, as the decoder effectively acts as a generator to enrich fine-grained acoustic detail.
In contrast, the semantic-only reconstruction lacks the speaker-specific fundamental frequencies and formants, yet retains clear semantic content. This demonstrates SAC’s ability to effectively separate and preserve the semantic features, while discarding speaker-dependent acoustic information. On the other hand, the acoustic-only reconstruction retains well-defined fundamental frequencies and formants, but completely loses semantic content (The output is perceptually similar to noise, devoid of any meaningful information). This highlights SAC’s remarkable capacity to disentangle and independently process semantic and acoustic features, providing direct evidence of its robust disentanglement ability. Additional mel-spectrogram reconstruction analyses of other codec models are provided in Appendix I.
@@ 6 Conclusion
In this paper, we presented SAC, a neural speech codec with semantic–acoustic dual-stream quantization. To exploit the strengths of semantic tokenizers for capturing linguistic content and codecs for modeling acoustic detail, SAC introduces independent semantic and acoustic streams that extract tokens separately. To further enhance timbre modeling, we incorporated speaker feature supervision into codec training. Comprehensive evaluations demonstrate that SAC achieves SOTA performance in both speech reconstruction and semantic representation across different bitrates, while ablation studies validate the effectiveness of auxiliary feature supervision. Moreover, downstream LLM-based TTS experiments further confirm SAC’s effectiveness as a speech tokenizer for generative speech applications. We also observe a remarkably clean disentanglement between semantic and acoustic tokens in reconstruction: semantic-token-based reconstruction contains no speaker-related information, while acoustic-token-based reconstruction preserves no semantic content. To the best of our knowledge, this is the first instance of such a clean disentanglement between semantics and speaker identity in terms of reconstruction. This finding offers new insights for subsequent tasks that rely on disentanglement, such as voice conversion or the joint control of style and timbre in TTS.
@@
Although SAC demonstrates superior reconstruction quality and semantic representation within the speech domain, its generalizability to other types of audio signals, such as music and sounds, remains to be explored. A key challenge lies in the semantic tokenizer used in SAC, which is trained under ASR supervision on speech data and therefore primarily aligned with textual objectives. In contrast, the semantics in music and sound extend beyond linguistic alignment. To develop a more general-purpose audio codec based on the SAC dual-stream architecture, future work should focus on designing a universal audio semantic encoder, potentially through multi-task supervision across modalities or self-supervised training on diverse audio data.
@@ Appendix A Details of Training Data
As summarized in Table 6, SAC is trained on approximately 20,000 hours of bilingual (Chinese and English) speech, comprising around 8 million utterances. The training corpus includes roughly 10,000 hours in each language. All audio is resampled to 16 kHz.
For Chinese, the data sources include the Chinese subset of Emilia (emilia), WenetSpeech4TTS (wenetspeech4tts), and a small amount of in-house para-linguistic data. For English, we use LibriSpeech (librispeech), the small and medium subsets of Libriheavy (libriheavy), the English subset of Emilia, MLS (mls), and a small amount of in-house data.
To promote diversity, we follow the annotation scheme of VoxBox (spark-tts), which labels speech samples along three dimensions: age, gender, and emotion. Specifically, we partition Emilia, WenetSpeech4TTS, and MLS into categories defined by the unique combination of these three attributes, and then sample data to achieve balanced coverage across categories. This strategy ensures broad diversity within the training set.
@@ Appendix B Codec Training Details
To accelerate SAC training, we pre-extracted all semantic representations $\mathbf{S}_{c}$ and corresponding semantic tokens from the training corpus using a pretrained semantic tokenizer (glm4voice-tokenizer). During training, only the codebook from the tokenizer is required to obtain the quantized semantic embeddings $\mathbf{S}_{q}$, while $\mathbf{S}_{c}$ serves as the ground-truth target for auxiliary semantic feature supervision. The generator contains approximately 277M parameters, with around 249M being trainable. The codebook of the semantic tokenizer and the speaker encoder are frozen to ensure that well-pretrained semantic and speaker features are properly utilized.
For generator training, the loss coefficients are set as $\lambda_{\text{recon}}=15$, $\lambda_{\text{vq}}=1$, $\lambda_{\text{adv}}=1$, $\lambda_{\text{feat}}=2$, $\lambda_{\text{sem}}=1000$, and $\lambda_{\text{spk}}=10$. Within the VQ loss, the commitment and codebook loss weights are set to 0.25 and 4, respectively. To improve training stability, the generator is pretrained for 1,500 steps before introducing the discriminator for adversarial training. An exponential moving average (EMA) is applied to maintain a smoothed version of the model parameters (AUV), which are used during inference and observed to enhance model stability.
@@ Appendix C Speech Reconstruction under Noisy Conditions
Notably, SAC maintains a clear advantage in UTMOS, SIM, and WER. For objective naturalness, SAC achieves UTMOS scores of 3.84 and 3.90 at high and low bitrates, respectively—both substantially higher than the ground truth value of 3.50. We attribute this to SAC’s acoustic stream effectively modeling fine-grained acoustic details, while the decoder functions as a generator that enhances objective perceptual realism. In terms of speaker similarity, SAC shows only a marginal drop of 0.01 in SIM compared to the clean condition, whereas other codecs experience significant declines. This demonstrates the robustness of both our model architecture and the diverse training data, as well as the effectiveness of the speaker feature supervision in preserving timbre characteristics. Furthermore, SAC maintains high speech intelligibility under noisy conditions, with remarkably low WER values, confirming that its dual-stream architecture enables reliable semantic reconstruction even in the presence of noise.
@@
Appendix D Downstream LLM-Based Speech Generation
To validate the potential of SAC in downstream speech language models (SLMs) tasks, we further conducted an LLM-based text-to-speech (TTS) experiment. For language modeling, we adopt a pure autoregressive (AR) framework using the pre-trained LLM Qwen3-0.6B  [^2] https://huggingface.co/Qwen/Qwen3-0.6B (qwen3) as the backbone. In contrast to prior TTS systems such as VALL-E (valle), which rely on a two-stage AR+NAR generation pipeline, our approach uses a single Transformer decoder to autoregressively predict single-layer speech tokens, substantially simplifying the modeling process.
To accommodate SAC’s dual-stream tokens, we introduce a novel interleaved flattening strategy for language modeling. Since SAC produces single-layer semantic and acoustic tokens with different token rates, we arrange them in an interleaved sequence proportional to their rate ratio. For example, the 37.5 Hz SAC variant generates semantic tokens at 12.5 Hz and acoustic tokens at 25 Hz; the downstream model therefore predicts them in a fixed 1:2 pattern in a single layer, where earlier semantic predictions facilitate subsequent acoustic predictions.
Given computational constraints, we use the 37.5 Hz SAC tokenizer and expand the LLM vocabulary with both semantic and acoustic codebooks. During training, the decoder-only LLM is optimized via negative log-likelihood to predict speech tokens conditioned on text transcriptions as prefixes.
Small-scale TTS evaluations often fail to reflect the robustness and generalization abilities of speech tokens, especially when paired with LLMs whose interaction dynamics exhibit complex emergent behavior. To obtain a more reliable assessment, we train the TTS model using SAC tokens on a large 100k-hour bilingual corpus (Chinese and English), following the same data distribution as VoxBox (spark-tts).
We evaluate the zero-shot TTS performance on the Seed-TTS-eval  (seed-tts), using speaker similarity (SIM), WER (or CER for the Chinese test set), and UTMOS as evaluation metrics. During inference, we concatenate the prompt text, the target synthesis text, and the prompt-audio token sequence as the input to the LLM. For a fair comparison, we benchmark our model against two state-of-the-art purely AR TTS systems trained on large-scale datasets: Spark-TTS (spark-tts) and Llasa (llasa-xcodec2). Spark-TTS adopts BiCodec as the speech tokenizer and Qwen2.5-0.5B as the backbone, trained on the same VoxBox data. Llasa employs X-codec2 as the tokenizer and Llama 3.2-1B (Llama) as the backbone, with versions trained on 80k, 160k, and 250k hours of data.
[…]
Seed-TTS test-en
Llasa-1B-80k 3.71 0.54 4.06
Llasa-1B-160k 3.60 0.56 4.05
Llasa-1B-250k 2.99 0.57 4.07
Spark-TTS 1.98 0.58 3.94
Ours 1.06 0.54 4.21
Seed-TTS test-zh
Llasa-1B-80k 2.69 0.65 3.27
Llasa-1B-160k 2.22 0.66 3.28
Llasa-1B-250k 1.89 0.67 3.28
Spark-TTS 1.20 0.67 3.27
Ours 0.90 0.65 3.34
Table 9: Results of our TTS model compared with prior single-stage AR TTS systems on the Seed-TTS test sets. The evaluation results for Spark-TTS and Llasa are cited from their respective papers.
Table 9 reports the zero-shot TTS performance of our model trained with the 37.5 Hz SAC tokenizer. Compared with prior pure AR TTS systems, our model achieves substantial gains in both intelligibility and objective naturalness. On the test-en set, our model attains a WER of 1.06%, markedly outperforming the Spark-TTS (1.98%). On the test-zh set, we obtain a CER of 0.90%, achieving a performance below 1% for the first time. To the best of our knowledge, our TTS model is optimal or near-optimal in terms of semantic clarity, even when compared against all existing TTS models (including non-purely AR models). Moreover, our model yields consistently higher UTMOS scores than both Llasa and Spark-TTS, demonstrating superior objective naturalness. These results collectively highlight the promising potential of SAC in downstream generative tasks and further confirm its capacity for high-fidelity semantic preservation and high-naturalness speech synthesis.
Despite these improvements, our model shows a slight reduction in speaker similarity compared with previous AR systems. We attribute this mainly to the low token rate of the 37.5 Hz SAC tokenizer: its coarse temporal granularity limits the capacity to capture fine-grained timbre details relative to the 50 Hz codecs used in prior TTS models.
Nevertheless, we believe that SAC has already demonstrated strong potential for generative downstream tasks. Moreover, for applications involving semantic or acoustic editing, the semantic–acoustic decoupled token design offers a particularly promising modeling foundation. As future work, we plan to investigate how data scaling and model scaling influence generation quality—especially speaker similarity—and to evaluate higher-rate SAC variants (e.g., 62.5 Hz) in zero-shot TTS, which remains unexplored due to current computational limitations.
Appendix E Subjective Evaluation on Speech Reconstruction
Since reconstructed audio from different codec models at high bitrates (particularly above 800 bps) tends to be perceptually indistinguishable to human listeners, we focus our subjective evaluation on low-bitrate settings. In this regime, Encodec (encodec) and SpeechTokenizer (speechtokenizer) produce notably low reconstruction quality when restricted to their first-layer RVQ tokens (with STOI scores below 0.80), and TS3-Codec (Ts3-codec) does not provide publicly available model weights. Consequently, these models are excluded from our comparison. Our Mean Opinion Score (MOS) study therefore includes Ground Truth, SemanticCodec (semanticodec), X-codec (xcodec), WavTokenizer (wavtokenizer), and SAC (with a token rate of 37.5 Hz).
For the evaluation, 20 native speakers were invited to assess reconstructed audio samples generated by each model. A total of 30 utterances were randomly selected from the LibriSpeech test-clean set. Evaluators were thoroughly informed of the scoring criteria and instructed to judge naturalness and perceptual quality as the primary factors in their evaluation. Each sample was rated on a 1–5 scale with 0.5-point increments, where higher scores indicate better perceived quality.
[…]
Ground Truth     - - 3.98
SemantiCodec     50 700 2.80
X-codec     50 500 3.18
WavTokenizer     40 480 3.20
SAC (ours)     37.5 525 3.94
Table 10: Subjective evaluation of reconstructed audio from different codec models on low-bitrate settings.
As shown in Table 10, SAC achieves a MOS of 3.94, substantially higher than all other codec baselines and approaching the ground-truth score of 3.98. This indicates that SAC preserves perceptual quality with almost no audible degradation, whereas competing codecs introduce noticeable distortion. The subjective results also highlight the limitations of UTMOS as a model-based predictor: although it reliably captures overall quality trends, finer-grained distinctions still require human evaluation. For example, WavTokenizer receives a lower UTMOS score than X-Codec (3.57 vs. 3.84), yet is slightly preferred in the human MOS assessment.
Appendix F Impact of Training Data Scale
Existing codec models are often trained on datasets of varying scales, which complicates a fair comparison of their performance. For example, WavTokenizer (wavtokenizer) is trained on 8k hours of mixed data, XY-Tokenizer (XY-Tokenizer) utilizes 101k hours of the multilingual Emilia (emilia) dataset, and TS3-Codec (Ts3-codec) uses 60k hours of Libri-light (libri-light) data. To validate the generality and scalability of the SAC modeling approach, we further include results trained exclusively on the LibriSpeech (librispeech) 960-hour English speech corpus (based on the low-bitrate version of SAC). This represents the smallest training data volume among all comparable bitrate codecs.
[…]
PESQ
[…]
PESQ
[…]
$\text{SAC}_{large}$     0.90 2.74 2.18 4.27 0.78 2.53
$\text{SAC}_{small}$     0.90 2.70 2.16 4.30 0.63 2.58
Table 11: Reconstruction evaluation of SAC trained with different data scales. $\text{SAC}_{large}$ denotes the model trained on the 20k hours of large-scale data mentioned in the main paper, while $\text{SAC}_{small}$ represents the model trained exclusively on the LibriSpeech dataset.
As shown in Table 11, $\text{SAC}_{small}$, trained on a small-scale dataset, still maintains superior reconstruction performance. Although its performance slightly declines in metrics such as PESQ and WER compared to $\text{SAC}_{large}$, it still demonstrates a significant advantage over other comparable-bitrate codecs presented in Table 2. Notably, $\text{SAC}_{small}$ achieves a further improvement in UTMOS, reaching up to 4.30. We primarily attribute this to the high purity of the audio in the LibriSpeech dataset, which enables the model to achieve better reconstruction quality on the equally clean test set (LibriSpeech test-clean) than $\text{SAC}_{large}$.
However, $\text{SAC}_{small}$ exhibits a substantial drop in SIM compared to $\text{SAC}_{large}$, decreasing from 0.78 to 0.63. We hypothesize that this is primarily caused by the severe lack of speaker diversity in the training data, leading SAC to overfit to speaker-specific characteristics. Specifically, we randomly sampled 2620 samples—the same number as the LibriSpeech test-clean set—from the training sets of both models and tested their reconstruction performance. The results show that the SIM score of $\text{SAC}_{small}$ on its training examples reached 0.72, which is significantly higher than 0.63 on the test set. In contrast, the SIM score of $\text{SAC}_{large}$ on its training examples remained consistent with the test examples (both 0.78), fully illustrating the robustness and generalization capability provided by large-scale training data.
In summary, this analysis indicates that SAC can maintain outstanding reconstruction performance even in low-resource scenarios, demonstrating excellent semantic clarity and acoustic quality. Furthermore, scaling up data, particularly by increasing the speaker diversity of the training corpus, effectively mitigates the issue of model overfitting to speaker characteristics, which further validates the generalization and scalability of the SAC model.
Appendix G Speech Reconstruction Speed
To evaluate the real-time performance of different speech codec models in speech reconstruction, we conduct experiments that measure the Real-Time Factor (RTF), defined as the ratio of processing time to audio duration. Specifically, 1,000 samples are randomly selected from the seed-eval dataset (seed-tts), and reconstruction is performed on an NVIDIA L20 GPU with a batch size of 1. For comparison, SAC is evaluated alongside two speech codecs of similar model scale—SemantiCodec (semanticodec) and XY-Tokenizer (XY-Tokenizer).
[…]
SemantiCodec    1400 507M 0.3608
XY-Tokenizer    1000 520M 0.0155
SAC    875 533M 0.0135
Table 12: Real-Time Factor (RTF) for audio codec models on test audio clips using an L20 GPU.
As shown in Figure 12, despite containing over 500M parameters, SAC maintains superior real-time performance compared to codec models of similar size, achieving an RTF as low as 0.0135. This efficiency primarily stems from SAC’s architecture: apart from the semantic tokenizer, which employs Transformer blocks, all remaining components consist of lightweight convolutional and linear layers that significantly accelerate inference. These results demonstrate SAC’s high efficiency and practicality for real-world deployment and various downstream applications.
Appendix H Effect of Auxiliary Feature Supervision on Semantic Representation
To investigate the impact of semantic and speaker feature supervision on SAC’s semantic representation capabilities, we further evaluate the semantic representation performance of SAC on ARCH under these two ablation settings.
[…]
SAC     61.81 39.63 29.21 99.63 57.57
   w/o $\mathcal{L}_{\text{spk}}$     59.72 40.32 27.34 99.54 56.73
   w/o $\mathcal{L}_{\text{sem}}$     62.19 38.78 29.14 99.60 57.43
Table 13: Ablation study of auxiliary feature supervision on semantic representations evaluations.
As shown in Table 13, removing the speaker feature supervision loss ($\mathcal{L}_{\text{spk}}$) or the semantic feature supervision loss ($\mathcal{L}_{\text{sem}}$) during training has a minimal impact on the final semantic representation evaluation results. Although the results fluctuate slightly across individual sub-evaluation sets, the overall score of the ablation models shows a decline of less than 1% compared to the baseline. This suggests that the auxiliary supervision branches have a limited effect on the semantic representation capabilities learned by SAC, and the dual-stream quantization design itself is sufficient to ensure the model retains rich semantic information.
Nevertheless, to achieve optimal semantic representation and reconstruction performance, we retained both feature supervision mechanisms in the final model design. However, for training efficiency, we also recommend removing the semantic supervision branch in resource-constrained scenarios, as this simplification only leads to a marginal decrease in both reconstruction performance and semantic representation capability.
Appendix I Mel-Spectrogram Reconstructions Comparison among Codecs